OpenStudy (anonymous):

Is the sequence geometric? 2, 6, 18, 54,... A. yes B. no

OpenStudy (mertsj):

A

OpenStudy (amistre64):

if 6/2 = 54/18 then yes

OpenStudy (amistre64):

otherwise no

OpenStudy (mertsj):

You always multiply by 3 to get the next term

OpenStudy (anonymous):

Okay thanks! I am writing that in my notebook so I do not forget that! :)

OpenStudy (amistre64):

the multiply by 3 part, or the dividing last terms part?

OpenStudy (anonymous):

Both!

OpenStudy (amistre64):

given a sequence\[a_1,a_2,a_3,a_4,...a_n,a_{n+1}\]we can determine if its geometric by taking the ratio any 2 terms that are next to each other.\[\frac{a_2}{a_1}=\frac{a_7}{a_6}=\frac{a_{n+1}}{a_n}\]

OpenStudy (amistre64):

of course it should equal if you go the other way to\[\frac{a_1}{a_2}=\frac{a_n}{a_{n+1}}\]but it just wont be the correct ratio